OpenStudy (anonymous):

A street that is 274 m long is covered in snow. City workers are using a snowplow to clear the street. The snowplow has tires that are 1.9 m in diameter. How many times does a tire have to turn in traveling the length of the street? Use the value 3.14 for pi . Round your answer to the nearest tenth. Do not round any intermediate steps.

OpenStudy (anonymous):

\[\frac{274}{dπ} = \frac{274}{1.9π} = ?\]

OpenStudy (ganpat):

Perimeter for tyre/ circle = 2* pi * r = 2 * 3.14 * 1.9 = 11.932 In one run, tier covers this distance.. so total number of turns = 274/ 11.932 = 22.9 = 23 turns..

OpenStudy (ganpat):

is that in ur options ??